# An Exploratory Study of V-Model in Building ML-Enabled Software: A Systems Engineering Perspective

### v3

@@ Abstract.
Machine learning (ML) components are being added to more and more critical and impactful software systems, but the software development process of real-world production systems from prototyped ML models remains challenging with additional complexity and interdisciplinary collaboration challenges. This poses difficulties in using traditional software lifecycle models such as waterfall, spiral, or agile models when building ML-enabled systems. In this research, we apply a Systems Engineering lens to investigate the use of V-Model in addressing the interdisciplinary collaboration challenges when building ML-enabled systems. By interviewing practitioners from software companies, we established a set of 8 propositions for using V-Model to manage interdisciplinary collaborations when building products with ML components. Based on the propositions, we found that despite requiring additional efforts, the characteristics of V-Model align effectively with several collaboration challenges encountered by practitioners when building ML-enabled systems. We recommend future research to investigate new process models that leverage the characteristics of V-Model such as the system decomposition, clear system boundary, and consistency of Validation & Verification (V&V) for building ML-enabled systems.
@@ 2.1. Software Engineering with ML Components
From the expertise perspective, ML-enabled systems impose expertise in a number of domains including data science, data quality (Sambasivan et al., 2021; Møller et al., 2020; Sambasivan and Veeraraghavan, 2022; Schelter et al., 2018), testing (Zhang et al., 2020; Riccio et al., 2020; Braiek and Khomh, 2020; Huang et al., 2020), security & privacy (Huang et al., 2011; Liu et al., 2018; McGraw et al., 2020; Wilhjelm and Younis, 2020), safety (Borg et al., 2019; Salay et al., 2017; Salay and Czarnecki, 2018; Zendel et al., 2015). These expertises come from different roles with different backgrounds and skill sets. Due to the introduction of ML components, these different roles, such as software engineer, ML scientist, technical leader, need to collaborate, and there are very few studies to address the challenges arising in these types of collaboration (Nahar et al., 2022). Regarding the lifecycle models for ML-enabled systems, ML lifecycle models such as CRISP-DM (Martínez-Plumed et al., 2019) and Team Data Science Process (TDSP) (Ericson et al., 2017) were studied in the Fintech domain, with findings that ML lifecycle models need to be revised with more research. In particular, these ML lifecycle models missed several necessary steps, including documentation, model evaluation, model monitoring, etc. Study that explored how teams develop, deploy and maintain software systems with ML components was conducted and summarized into best practices for improving software development processes  (Serban et al., 2020). In this work, we attempt to address the collaboration problem from the perspective of a more interdisciplinary-focused process model.
@@ 4.1. Vee Chart
SE1
Mission Objectives and Constraints
Objectives
Statements describing the goals and constraints of the (mission) objectives
SE2
Derived Requirements Development
Derived Requirements Development
Concise statements outlining 1) what to achieve, 2) how the desired level of achievement, and 3) any associated constraints or limitations.
SE3
Architectural Design Development
Architectural Design Development
Explanation of elements and interfaces, with detailed analysis for assessing the performance
SE4
Concept of Operation (“ConOps”)
User Scenarios and Rationale
Explanation of how the system operates under the (mission) objectives to align with stakeholder expectations
SE5
Validate and Verify
Validate and Verify
An ongoing SE function, ensuring consistency in requirements, ConOps, and architectural design during formulation, while also planning and executing tests under a verification plan
SE6
Interfaces and ICD (Interface Control Document)
Documented Interfaces
Descriptions and documentations of interfaces, boundaries between elements, that evolve as the architectural design with requirements outlined.
SE7
Mission Environment
External Events
Descriptions and documentations of external environmental concerns and exposures
SE8
Technical Resource Budget Tracking
Resource Tracking
Tracking and identification of resource budgets
SE9
Risk Management
Risk Management
Negative occurrences and their potential unfavorable impacts on the project
SE10
Configuration Management and Documentation
Documentation and Approval
A system for documentation control, access, approval and dissemination
SE11
System Milestone Reviews and Reports
Milestone Reviews and Reports
Reviews that communicate to stakeholders with significant project events, decision points, and achievements, etc.
@@ 6. Conclusions, Limitations, and Future Work
In this work, we take a Systems Engineering lens to close the gap in collaboration challenges when building ML-enabled systems through the application of the V-Model. With the V-Model and the interviews with industrial practitioners, we collected empirical evidence and then developed 8 propositions to address the research question of how can software process life cycles be established to address the interdisciplinary collaboration challenges from ML-enabled systems? Our study revealed that, despite requiring additional efforts, the characteristics of V-Model, such as the system decomposition, clear boundaries and responsibilities between systems and components, and consistency of V&V, align effectively with several collaboration challenges encountered by practitioners when building ML-enabled systems. This work makes contributions by 1) conducting one of the first exploratory studies to apply V-Model in building ML-enabled systems, and 2) addressing the gap of these collaboration challenges with a set of 8 propositions and their connections to V-Model. We recommend future research to investigate new process models that leverage these characteristics of V-Model for building ML-enabled systems.

### v4

@@ Abstract.
Machine learning (ML) components are being added to more and more critical and impactful software systems, but the software development process of real-world production systems from prototyped ML models remains challenging with additional complexity and interdisciplinary collaboration challenges. This poses difficulties in using traditional software lifecycle models such as waterfall, spiral, or agile models when building ML-enabled systems. In this research, we apply a Systems Engineering lens to investigate the use of V-Model in addressing the interdisciplinary collaboration challenges when building ML-enabled systems. By interviewing practitioners from software companies, we established a set of 8 propositions for using V-Model to manage interdisciplinary collaborations when building products with ML components. Based on the propositions, we found that despite requiring additional efforts, the characteristics of V-Model align effectively with several collaboration challenges encountered by practitioners when building ML-enabled systems. We recommend future research to investigate new process models, frameworks and tools that leverage the characteristics of V-Model such as the system decomposition, clear system boundary, and consistency of Validation & Verification (V&V) for building ML-enabled systems.
@@ 2.1. Software Engineering with ML Components
From the expertise perspective, ML-enabled systems impose expertise in a number of domains including data science, data quality (Sambasivan et al., 2021; Møller et al., 2020; Sambasivan and Veeraraghavan, 2022; Schelter et al., 2018), testing (Zhang et al., 2020; Riccio et al., 2020; Braiek and Khomh, 2020; Huang et al., 2020), security & privacy (Huang et al., 2011; Liu et al., 2018; McGraw et al., 2020; Wilhjelm and Younis, 2020), safety (Borg et al., 2019; Salay et al., 2017; Salay and Czarnecki, 2018; Zendel et al., 2015). These expertises come from different roles with different backgrounds and skill sets. Due to the introduction of ML components, these different roles, such as software engineer, ML scientist, technical leader, need to collaborate, and there are very few studies to address the challenges arising in these types of collaboration (Nahar et al., 2022). Regarding the lifecycle models for ML-enabled systems, ML lifecycle models such as CRISP-DM (Martínez-Plumed et al., 2019) and Team Data Science Process (TDSP) (Ericson et al., 2017) were studied in the Fintech domain, with findings that ML lifecycle models need to be revised with more research (Haakman et al., 2021). In particular, these ML lifecycle models missed several necessary steps, including documentation, model evaluation, model monitoring, etc. Study that explored how teams develop, deploy and maintain software systems with ML components was conducted and summarized into best practices for improving software development processes  (Serban et al., 2020). In this work, we attempt to address the collaboration problem from the perspective of a more interdisciplinary-focused process model.
@@ 4.1. Vee Chart
SE1 Mission Objectives and Constraints Objectives Statements describing the goals and constraints of the (mission) objectives
SE2 Derived Requirements Development Derived Requirements Development Concise statements outlining 1) what to achieve, 2) how the desired level of achievement, and 3) any associated constraints or limitations.
SE3 Architectural Design Development Architectural Design Development Explanation of elements and interfaces, with detailed analysis for assessing the performance
SE4 Concept of Operation (“ConOps”) User Scenarios and Rationale Explanation of how the system operates under the (mission) objectives to align with stakeholder expectations
SE5 Validate and Verify Validate and Verify An ongoing SE function, ensuring consistency in requirements, ConOps, and architectural design during formulation, while also planning and executing tests under a verification plan
SE6 Interfaces and ICD (Interface Control Document) Documented Interfaces Descriptions and documentations of interfaces, boundaries between elements, that evolve as the architectural design with requirements outlined.
SE7 Mission Environment External Events Descriptions and documentations of external environmental concerns and exposures
SE8 Technical Resource Budget Tracking Resource Tracking Tracking and identification of resource budgets
SE9 Risk Management Risk Management Negative occurrences and their potential unfavorable impacts on the project
SE10 Configuration Management and Documentation Documentation and Approval A system for documentation control, access, approval and dissemination
SE11 System Milestone Reviews and Reports Milestone Reviews and Reports Reviews that communicate to stakeholders with significant project events, decision points, and achievements, etc.
@@ 6. Conclusions, Limitations, and Future Work
In this work, we take a Systems Engineering lens to close the gap in collaboration challenges when building ML-enabled systems through the application of the V-Model. With the V-Model and the interviews with industrial practitioners, we collected empirical evidence and then developed 8 propositions to address the research question of how can software process life cycles be established to address the interdisciplinary collaboration challenges from ML-enabled systems? Our study revealed that, despite requiring additional efforts, the characteristics of V-Model, such as the system decomposition, clear boundaries and responsibilities between systems and components, and consistency of V&V, align effectively with several collaboration challenges encountered by practitioners when building ML-enabled systems. This work makes contributions by 1) conducting one of the first exploratory studies to apply V-Model in building ML-enabled systems, and 2) addressing the gap of these collaboration challenges with a set of 8 propositions and their connections to V-Model. We recommend future research to investigate new process models, frameworks and tools that leverage these characteristics of V-Model for building ML-enabled systems.